# CryoNet.Refine: A One-step Diffusion Model for Rapid Refinement of Structural Models with Cryo-EM Density Map Restraints

### v1

@@ 1 Introduction
The landscape of AI in structural biology has been dramatically reshaped by recent advancements in generative models, particularly diffusion models, which excel in protein generation, design, and structure prediction. Breakthroughs like AlphaFold3(Abramson et al., 2024), RFDiffusion(Watson et al., 2023), and Chroma(Ingraham et al., 2023) exemplify their exceptional ability to generate diverse and high-quality structures for various biomolecules, including proteins, DNA, and RNA complexes. Beyond single-chain protein prediction, where models like AlphaFold2(Jumper et al., 2021), RoseTTAFold (Baek et al., 2021), and ESMFold(Lin et al., 2023) have achieved remarkable accuracy, diffusion-based methods are extending to protein-protein interactions, protein-small molecule complexes, and RNA/DNA complexes (Baek et al., 2021; 2023; Wohlwend et al., 2025; Passaro et al., 2025). Notably, AlphaFold3 has moved beyond predefined fixed bond lengths and angles from AlphaFold2, by learning these geometric constraints directly from PDB data using diffusion models. This includes accurately capturing geometry features like the planarity of benzene rings in amino acids. However, a significant limitation of these generative methods is that while they produce geometrically plausible structures, they often exhibit suboptimal performance on various detailed geometric metrics and, crucially, do not natively support refinement under the direct restraints of experimental data, such as cryo-EM density maps. Applying the robust generative ability of diffusion models to cryo-EM data-restraints refinement thus offers a transformative pathway to substantially improve structure quality and overcome the reliance on laborious, parameter-heavy traditional or manual refinement workflows.
@@
To address these limitations, we introduce CryoNet.Refine, a novel one-step diffusion model designed for cryo-EM atomic model refinement Figure 1.(a). CryoNet.Refine pioneers the integration of advanced atomic generation principles, inspired by the capabilities seen in AlphaFold3, into a comprehensive refinement framework. Taking a cryo-EM density map and an atomic structure as input, Atom encoder first extracts intricate features from the structure to be refined. Concurrently, Sequence embedder encodes atomic type information derived from the input molecular sequence. These encoded representations are then fed into our one-step diffusion module, which iteratively generates the refined atomic structure. The refinement process is guided by novel designed density loss and a set of geometry losses, facilitating iterative optimization of the atomic model. CryoNet.Refine achieves the first successful atomic structure refinement under the direct constraints of experimental cryo-EM density maps and the guidance of standard geometry metrics within a neural network framework (Figure 1.(b)). Experimental benchmarkds demonstrates significant improvements over Phenix.real_space_refine on both protein-protein and DNA/RNA-protein complexes, showcasing its exceptional performance and broad applicability. Our work presents four main contributions:
@@ 3 Methods
CryoNet.Refine is an end-to-end deep learning framework designed for the atomic model refinement of macromolecules directly against experimental cryo-EM density maps Figure 2.(a). The refinement process begins by taking an experimental cryo-EM density map (${d}_{0}$) and an initial atomic structure (${x}_{0}$) as input. Atom encoder first processes the input structure to extract pairwise features (z{z}) for each atom. Concurrently, Sequence embedder encodes atomic type information (s{s}) derived from the input molecular sequence. These encoded representations (z{z},s{s}) along with the initial atomic structure (${x}_{0}$) are then fed into our one-step Diffusion Module to generate a refined atomic structure (${x}_{1}$). Subsequently, the density generator takes the refined structure (${x}_{1}$) as input to produce a simulated density map (${d}_{1}$). A designed density loss ($\mathcal{L}_{\text{den}}$) is then calculated by comparing this simulated map (${d}_{1}$) with the input map (${d}_{0}$). Simultaneously, a Geometry loss ($\mathcal{L}_{\text{geo}}$) is computed based on the refined structure (${x}_{1}$). These two losses are then weighted and summed to form the total loss ($\mathcal{L}_{\text{total}}$), which is backpropagated to update the parameters of the diffusion module. This process constitutes one refinement cycle. The refined structure from this cycle (${x}_{1}$) is then used as the input for the next recycle step to generate ${x}_{2}$. After n{n} such cycles, the final refined atomic model, ${x}_{n}$, is obtained. Our network was first initialized with the pre-trained parameters of Boltz-2(Passaro et al., 2025), a PyTorch implementation of AlphaFold3, followed by training with a composite loss function including two components: one measuring experimental data fidelity and the other introducing geometric restraints. The network was trained until convergence, yielding the generated atomic structures that are in high agreement with the experimental data while satisfying geometric constraints.
@@
Figure 2: (a) Overview of the CryoNet.Refine framework. The CryoNet.Refine Network consists of four modules: Atom encoder, Sequence embedder, Diffusion module, and Density generator. The input atomic model is first processed by the encoders, and the resulting features are fed into a one-step diffusion module to generate an initial refined atomic model. Subsequently, the Density generator creates a synthetic density map, which is used to compute density loss against the input density map and geometry loss based on geometry restraints. These losses are then backpropagated to optimize the diffusion module, while the atomic model is further refined through multiple recycle steps until convergence. (b) Model-map correlation coefficient trajectory over 234 recycling CryoNet.Refine iterations on the structure of the human concentrative nucleoside transporter CNT3((Bank, )). The input density map is EMD-0775, the input atomic model is predicted by AlphaFold3.
@@
where $c_{\text{skip}}$$c_{\text{out}}$ and $c_{\text{out}}$ denote coefficients in the preconditioned forward module(Appendix  C) and $\mathcal{F}_{\theta}$ is a parameterized neural network in it. $\mathcal{C}$ denotes conditioned features derived from encoded structural features $s$ and $z$. This preconditioned update ensures that the one-step refinement preserves the theoretical properties of multi-step diffusion while collapsing the stochastic process into a deterministic, single-step prediction.
Restraints of density map and structure geometry. CryoNet.Refine employs a density generator and two types of loss functions to leverage these restraints. The first type of loss is a density loss, derived by compute the correlation between the experimental cryo-EM maps and the synthetic maps generated from refined atomic model by the density generator (Section 3.2.1). Another one is a set of geometry loss terms imposes stereochemical restraints—such as those derived from Ramachandran plot, rotamer, and bond angle distributions—to ensure standard structure geometry (Section 3.2.2). Thus the network is trained using a recycling strategy that reuses refined outputs as inputs for subsequent passes, progressively optimizing the two loss functions for density map agreement and geometry metrics. Together, these novel design choices enable CryoNet.Refine to achieve rapid, generalizable, and experimentally consistent refinement by fully respecting the restraints derived from both the cryo-EM density map and the structure geometry.
@@ 3.2.1 Density loss
We compute the density loss within two steps: (i) generate a synthetic density map for the refined atomic model; (ii) compute the overlap region between the input density and the synthetic density. Both these two steps must be implemented to be fully differentiable.
A density map $d$ can be discretized as a group of 4-dimensional vectors ${\lambda},{\mu},{\nu},\rho]$, where $\vec{{\bm{m}}}=[{\lambda},{\mu},{\nu}]$ specifies a position and $\rho$ is the density value at this point, with the origin of coordinates nearest to $\vec{{\bm{m}}}_{\text{min}}$ and farthest to $\vec{{\bm{m}}}_{\text{max}}$. For computation of the overlapped region, we first locate this region $\forall\vec{{\bm{m}}}_{o}\in\mathcal{S}_{o}$ between the input density map $d_{0}$ and the synthetic density map $d_{i}$ through the coordinate system:
After locating $\mathcal{S}_{o}$, we can select all the density values within it from $d_{0}$ and $d_{i}$ and get their subsets $d_{0\bigcap o}$ and $d_{i\bigcap o}$. Thus, the density loss will be defined as:
[…]
For the implementation of synthetic density map generation, we follow the molmap algorithm in ChimeraX(Pettersen et al., 2021), utilizing the all-atom coordinates of the refined atomic model at the $i$ recycle $\vec{{\bm{x}}}_{i}=[\bm{x}_{i}^{0},\bm{x}_{i}^{1},\bm{x}_{i}^{2}]$ and output a box-shaped density map $d_{i}=(\vec{{\bm{m}}}_{i},\bm{\rho}_{i})$, where $\vec{{\bm{x}}}_{i}\in\mathbb{R}^{3\times N_{\text{atom}}}$ and $\vec{{\bm{m}}}_{i}\in\mathbb{R}^{3\times N_{\text{atom}}\times L}$, noting that the parameter $L$ stands for the number of grid points for each atom per axis, and its value is obtained by:
The key idea here is to construct a Gaussian sphere for each atom, with $r_{\mathcal{G}}$ and $s_{\mathcal{G}}$ respectively meaning Gaussian radius and the radius of a Gaussian sphere. We expect that $\vec{{\bm{m}}}_{i}$ turns out to be a box as small as possible with all-atom coordinates $\vec{{\bm{x}}}_{i}$ inside it as well as a margin of Gaussian sphere around it, so:
$\vec{m}_{i_{\text{min}}}=[\lfloor\min(\bm{x}_{i}^{0})\rfloor,\lfloor\min(\bm{x}_{i}^{1})\rfloor,\lfloor\min(\bm{x}_{i}^{2})\rfloor],\quad$ $\vec{m}_{i_{\text{max}}}=[\lceil\max(\bm{x}_{i}^{0})\rceil,\lceil\max(\bm{x}_{i}^{1})\rceil,\lceil\max(\bm{x}_{i}^{2})\rceil]$
$\vec{m}_{i\_\text{size}}=\lceil(\vec{m}_{i_{\text{max}}}-\vec{m}_{i_{\text{min}}})/v+2\cdot s_{\mathcal{G}}+1\rceil,\quad$ $\vec{m}_{i_{\text{center}}}=(\vec{m}_{i_{\text{min}}}+\vec{m}_{i_{\text{max}}})/(2\cdot v)$
$\vec{m}_{i_{\text{min}}}=\lfloor(\vec{m}_{i_{\text{center}}}-\vec{m}_{i\_\text{size}})/2\rfloor,\quad$ $\vec{m}_{i_{\text{max}}}=\lfloor(\vec{m}_{i_{\text{center}}}+\vec{m}_{i\_\text{size}})/2\rfloor$
$\text{for }n\in(1,2,...,N_{\text{atom}})\quad{\bm{\rho}}_{\lambda}^{n}=w_{n}\cdot\exp(-\frac{(x_{i}^{0})^{n}+{\bm{l}}}{2})\text{, }{\bm{l}}=[-s_{\mathcal{G}},...,0,-s_{\mathcal{G}}],$ (4)
@@ 3.2.2 Geometry loss
Geometry loss functions are employed to guarantee the stereochemical accuracy and structural validity of predicted proteins by enforcing conformity with established biological restraints. However, previous work has failed to incorporate geometry-related losses such as those for Ramachandran plot, rotamer, and C-beta deviation. Therefore, we innovatively implemented these differentiable geometry loss functions (ramachandran loss $\mathcal{L}_{\text{rama}}$, rotamer loss $\mathcal{L}_{\text{rot}}$, $C_{\beta}$ deviation loss) in this work.
@@ 4.1 Benchmark
We curated a benchmark of 63 complexes (53 protein, 10 DNA/RNA–protein) with cryo-EM density map resolutions ranging from 2.0 Åto 4.5 Å. For each case, initial atomic models were predicted by AlphaFold3 and subsequently refined against the experimental density map by Phenix.real_space_refine and CryoNet.Refine. Dataset statistics are provided in Figure 10 and Appendix B.
@@ 4.2 Refinement of Protein Complexes
Table 1: Performance on protein complex (↑: higher is better, ↓: lower is better).
@@
Model-map Correlation Coefficients $\rm CC_{mask}\uparrow$ 0.44 0.61 0.65
$\rm CC_{box}\uparrow$ 0.45 0.58 0.60
$\rm CC_{mc}\uparrow$ 0.45 0.62 0.65
$\rm CC_{sc}\uparrow$ 0.43 0.60 0.63
$\rm CC_{peaks}\uparrow$ 0.31 0.48 0.52
$\rm CC_{volume}\uparrow$ 0.47 0.61 0.65
@@
Figure 3 and Figure 4 show the performance of CryoNet.Refine across six model-map correlation coefficients (CC) and model geometric metrics, respectively. Notably, the largest gains in $\mathrm{CC_{mask}}$ and $\mathrm{CC_{mc}}$ suggest the refined atomic model more accurate main-chain placement within the density map. Regarding geometry, our method consistently achieves superior stereochemistry, virtually eliminating $C_{\beta}$ deviations, raising ramachandran favored to nearly 99%, and reducing rotamer outliers by over 55% compared to Phenix.real_space_refine. Furthermore, angle RMSD is drastically reduced from 1.56° to 0.48°. These results emphatically highlight that CryoNet.Refine simultaneously improve the stereochemical geometry and the agreement to the cryo-EM density map.
@@
Figure 4: Model geometric metrics. (Color gradient: blue for better, red for worse)
@@ 5.1 Runtime Performance
We benchmarked the runtime of CryoNet.Refine against Phenix.real_space_refine. Leveraging GPU parallelization, CryoNet.Refine consistently achieves highly efficient performance, whereas Phenix.real_space_refineś CPU-only support incurs higher computational costs, particularly for large complexes.
Across 63 complexes, CryoNet.Refine ran faster than Phenix.real_space_refine in 40 cases (63%). These results underscore that our method combines superior accuracy with high efficiency, making it ideal for large-scale, high-throughput cryo-EM refinement.

### v2

@@ 1 Introduction
The landscape of AI in structural biology has been dramatically reshaped by recent advancements in generative models, particularly diffusion models, which excel in protein generation, design, and structure prediction. Breakthroughs like AlphaFold3(Abramson et al., 2024), RFDiffusion(Watson et al., 2023), and Chroma(Ingraham et al., 2023) exemplify their exceptional ability to generate diverse and high-quality structures for various biomolecules, including proteins, DNA, and RNA complexes. Beyond single-chain protein prediction, where models like AlphaFold2(Jumper et al., 2021), RoseTTAFold (Baek et al., 2021), and ESMFold(Lin et al., 2023) have achieved remarkable accuracy, diffusion-based methods are extending to protein-protein interactions, protein-small molecule complexes, and DNA/RNA complexes (Baek et al., 2021; 2023; Wohlwend et al., 2025; Passaro et al., 2025). Notably, AlphaFold3 has moved beyond predefined fixed bond lengths and angles from AlphaFold2, by learning these geometric constraints directly from PDB data using diffusion models. This includes accurately capturing geometry features like the planarity of benzene rings in amino acids. However, a significant limitation of these generative methods is that while they produce geometrically plausible structures, they often exhibit suboptimal performance on various detailed geometric metrics and, crucially, do not natively support refinement under the direct restraints of experimental data, such as cryo-EM density maps. Applying the robust generative ability of diffusion models to cryo-EM data-restraints refinement thus offers a transformative pathway to substantially improve structure quality and overcome the reliance on laborious, parameter-heavy traditional or manual refinement workflows.
@@
To address these limitations, we introduce CryoNet.Refine, a novel one-step diffusion model designed for cryo-EM atomic model refinement Figure 1.(a). CryoNet.Refine pioneers the integration of advanced atomic generation principles, inspired by the capabilities seen in AlphaFold3, into a comprehensive refinement framework. Taking a cryo-EM density map and an atomic structure as input, Atom encoder first extracts intricate features from the structure to be refined. Concurrently, Sequence embedder encodes atomic type information derived from the input molecular sequence.The Pairformer module follows the Boltz-2(Passaro et al., 2025) perform cross-attention between the atom and the sequence embeddings. These representations are then fed into our one-step diffusion module, which iteratively generates the refined atomic structure. The refinement process is guided by novel designed density loss and a set of geometry losses, facilitating iterative optimization of the atomic model. CryoNet.Refine achieves the first successful atomic structure refinement under the direct constraints of experimental cryo-EM density maps and the guidance of standard geometry metrics within a neural network framework (Figure 1.(b)). Experimental benchmarkds demonstrates significant improvements over Phenix.real_space_refine on both protein-protein and DNA/RNA-protein complexes, showcasing its exceptional performance and broad applicability. Our work presents four main contributions:
@@ 3 Methods
CryoNet.Refine is an end-to-end deep learning framework designed for the atomic model refinement of macromolecules directly against experimental cryo-EM density maps Figure 2.(a). The refinement process begins by taking an experimental cryo-EM density map (${d}_{0}$) and an initial atomic structure (${x}_{0}$) as input. Atom encoder first processes the input structure to extract pairwise features (z{z}) for each atom. Concurrently, Sequence embedder encodes atomic type information (s{s}) derived from the input molecular sequence. These encoded representations (z{z},s{s}) along with the initial atomic structure (${x}_{0}$) are then fed into our one-step Diffusion Module to generate a refined atomic structure (${x}_{1}$). Subsequently, the density generator takes the refined structure (${x}_{i}$) as input to produce a simulated density map (${d}_{i}$). A designed density loss ($\mathcal{L}_{\text{den}}$) is then calculated by comparing this simulated map (${d}_{1}$) with the input map (${d}_{0}$). Simultaneously, a Geometry loss ($\mathcal{L}_{\text{geo}}$) is computed based on the refined structure (${x}_{i}$). These two losses are then weighted and summed to form the total loss ($\mathcal{L}$), which is backpropagated to update the parameters of the diffusion module. Crucially, this workflow represents a test-time optimization strategy rather than a static inference pass. For each specific case, the network undergoes this iterative training-and-optimization loop to fine-tune the parameters of the diffusion module, thereby customizing the atomic structure against the experimental density map and geometric constraints. This process constitutes one refinement cycle. After n{n} such cycles, the final refined atomic model, ${x}_{n}$, is obtained. Our network was first initialized with the parameters of Boltz-2(Passaro et al., 2025), a PyTorch implementation of AlphaFold3, followed by training with a composite loss function including two components: one measuring experimental data fidelity and the other introducing geometric restraints. The network was trained until convergence, yielding the generated atomic structures that are in high agreement with the experimental data while satisfying geometric constraints.
@@
Figure 2: (a) Overview of the CryoNet.Refine framework. The CryoNet.Refine Network consists of five modules: Atom encoder, Sequence embedder, Pairformer module, Diffusion module, and Density generator. The input atomic model is first processed by the encoders, and the resulting features are fed into a one-step diffusion module to generate an initial refined atomic model. Subsequently, the Density generator creates a synthetic density map, which is used to compute density loss against the input density map and geometry loss based on geometry restraints. These losses are then backpropagated to optimize the diffusion module, while the atomic model is further refined through multiple recycle steps until convergence. (b) Model-map correlation coefficient trajectory over 234 recycling CryoNet.Refine iterations on the structure of the human concentrative nucleoside transporter CNT3((Bank, )). The input density map is EMD-0775, the input atomic model is predicted by AlphaFold3.
@@
where $c_{\text{skip}}$, $c_{\text{out}}$ and $c_{\text{out}}$ denote coefficients in the preconditioned forward module(Appendix  C) and $\mathcal{F}_{\theta}$ is a parameterized neural network in it. $\mathcal{C}$ denotes conditioned features derived from encoded structural features $s$ and $z$. This preconditioned update ensures that the one-step refinement preserves the theoretical properties of multi-step diffusion while collapsing the stochastic process into a deterministic, single-step prediction.
Restraints of density map and structure geometry. CryoNet.Refine employs a density generator and two types of loss functions to leverage these restraints. The first type of loss is a density loss, derived by compute the correlation between the experimental cryo-EM map and the synthetic map generated from refined atomic model by the density generator (Section 3.2.1). Another one is a set of geometry loss terms imposes stereochemical restraints—such as those derived from Ramachandran plot, rotamer, and bond angle distributions—to ensure standard structure geometry (Section 3.2.2). Thus the network is trained using a recycling strategy that reuses refined outputs as inputs for subsequent passes, progressively optimizing the two loss functions for density map agreement and geometry metrics. Together, these novel design choices enable CryoNet.Refine to achieve rapid, generalizable, and experimentally consistent refinement by fully respecting the restraints derived from both the cryo-EM density map and the structure geometry.
@@ 3.2.1 Density loss
We compute density loss in two steps: (i) generate a synthetic density map for the refined atomic model; (ii) compute cosine similarity between the input and the synthetic density map to yield the loss term. Both these two steps must be implemented to be fully differentiable. Here we briefly describe how the differentiable density generator works and the formula for density loss function, with more details discussed in Appendix D.
Our current density generator is not a neural network but a physics-based simulator. Given the grid points $\bm{\vec{m}}$ and all-atom coordinates of the structure model $\vec{\mathbf{x}}$ with $N$ atoms, we get the density values $\hat{\bm{\rho}}$ by constructing Gaussian spheres centering at each atom position:
$\hat{\bm{\rho}}(\vec{{\bm{m}}},\vec{\mathbf{x}})=\sum_{i=1}^{N}w_{i}e^{-k|\vec{{\bm{m}}}-\vec{\mathbf{x}}_{i}|^{2}}$ (3)
where $w_{i}$ represents the atomic number of the $i^{th}$ atom, $\vec{\mathbf{x}}$ is the position vector of the $i^{th}$ atom, and $k$ is determined by the synthetic density map resolution resres as well as the voxel size $v$ (both specified as identical to the input experimental map):
$k=8\cdot{res}/(\pi\cdot v)$ (4)
Based on the synthetic density values $\hat{\bm{\rho}}(\vec{{\bm{m}}},\vec{\mathbf{x}})$ and those retrieved from the input density map $\bm{\rho}(\vec{{\bm{m}}})$, we can calculate the density loss function $\mathcal{L}_{\text{den}}$ in a fully differentiable manner:
$\mathcal{L}_{\text{den}}=1-\frac{\hat{\bm{\rho}}(\vec{{\bm{m}}},\vec{\mathbf{x}})\cdot\bm{\rho}(\vec{{\bm{m}}})}{||\hat{\bm{\rho}}(\vec{{\bm{m}}},\vec{\mathbf{x}}||\cdot||\bm{\rho}(\vec{{\bm{m}}})||}$ (5)
in which the second term is effectively the cosine similarity between $\hat{\bm{\rho}}(\vec{{\bm{m}}},\vec{\mathbf{x}})$ and $\bm{\rho}(\vec{{\bm{m}}})$. This Gaussian scattering formula to generate the synthetic density map conceptually follows the molmap method in UCSF ChimeraX(Pettersen et al., 2021), but the key difference is that we redesign the entire computation process to be fully differentiable using PyTorch. This fully differentiable implementation is essential for our method because it allows the density loss $\mathcal{L}_{\text{den}}$ to serve as a restraint term for back-propagation during each refinement loop. It’s also worth pointing out that our density generator achieves better performance than the UCSF ChimeraX(Pettersen et al., 2021) it molmap method, with the averaged correlation coefficients between synthetic and input density maps improved from 0.803 to 0.892.
@@ 3.2.2 Geometry loss
Geometry loss functions are employed to guarantee the stereochemical accuracy and structural validity of predicted proteins by enforcing conformity with established biological restraints. However, previous work has failed to incorporate geometry-related losses such as those for Ramachandran plot, rotamer, and Cβ deviation. Therefore, we innovatively implemented these differentiable geometry loss functions (ramachandran loss $\mathcal{L}_{\text{rama}}$, rotamer loss $\mathcal{L}_{\text{rot}}$, $C_{\beta}$ deviation loss) in this work.
@@ 4.1 Benchmark
We curated a benchmark of 120 complexes (110 protein, 10 DNA/RNA–protein) with cryo-EM density map resolutions ranging from 1.8 Åto 5.9 Å. For each case, initial atomic models were predicted by AlphaFold3 and subsequently refined against the experimental density map by Phenix.real_space_refine and CryoNet.Refine. Dataset statistics are provided in Figure 10 and Appendix B.
@@ 4.2 Refinement of Protein Complexes
Table 1: Performance on benchmark of 110 protein complexes(↑: higher is better, ↓: lower is better).
@@
Model-map Correlation Coefficients $\rm CC_{mask}\uparrow$ 0.38 0.54 0.59
$\rm CC_{box}\uparrow$ 0.41 0.53 0.57
$\rm CC_{mc}\uparrow$ 0.40 0.55 0.60
$\rm CC_{sc}\uparrow$ 0.39 0.55 0.58
$\rm CC_{peaks}\uparrow$ 0.27 0.40 0.45
$\rm CC_{volume}\uparrow$ 0.42 0.55 0.60
@@
Figure 3 and Figure 4 show the performance of CryoNet.Refine across six model-map correlation coefficients (CC) and model geometric metrics, respectively. Notably, the consistent gains in $\mathrm{CC_{mask}}$ (0.59 vs. 0.54) and $\mathrm{CC_{mc}}$ (0.60 vs. 0.55) suggest the refined atomic model achieves more accurate main-chain placement within the density map. Regarding geometry, our method consistently achieves superior stereochemistry, virtually eliminating $C_{\beta}$ deviations, raising Ramachandran favored percentage to nearly 99% (98.92%), and reducing rotamer outliers by approximately 57% compared to Phenix.real_space_refine (1.15% to 0.49%). Furthermore, Angle RMSD is drastically reduced from the initial 1.58 to 0.37. These results emphatically highlight that CryoNet.Refine simultaneously improves the stereochemical geometry and the agreement to the cryo-EM density map.
@@
Figure 4: Model geometric metrics. (Color gradient: blue for better, red for worse.)
@@ 5.1 Runtime Performance
We benchmarked the runtime of CryoNet.Refine against Phenix.real_space_refine. CryoNet.Refine consistently achieves highly efficient performance, whereas Phenix.real_space_refineś CPU-only support incurs higher computational costs, particularly for large complexes.
Across 120 complexes, CryoNet.Refine ran faster than Phenix.real_space_refine in 65 cases (54.2%). These results underscore that our method combines superior accuracy with high efficiency, making it ideal for large-scale, high-throughput cryo-EM refinement.